# Using Large Language Models for Template Detection from Security Event Logs

### v1

@@ IV-C Results
TABLE III: Performance of Template Detection Methods
Method Number of correctly detected templates Number of all detected templates F1-score
sshd su suricata snmpd apache sshd su suricata snmpd apache sshd su suricata snmpd apache
OpenChat 27 10 3 7 5 32 14 15 7 7 0.794 0.741 0.286 1 0.714
Mistral 17 9 - 5 1 25 9 - 6 3 0.557 0.818 - 0.769 0.2
Wizardlm2 9 1 - 5 3 50 9 - 6 7 0.209 0.091 - 0.769 0.429
Drain 29 9 3 6 2 44 16 80 8 4 0.725 0.621 0.07 0.8 0.364
@@
According to Table III, OpenChat achieved the best performance on four data sets, whereas Mistral had the highest F1-score on the su data set. The suricata data set was the most challenging for all evaluated approaches since Mistral and Wizardlm2 were not able to process it within 10 hours, while OpenChat and Drain achieved low F1-scores of 0.286 and 0.07, respectively. The vast majority of the log messages in this data set were Suricata IDS alerts with different numbers of words that correspond to one ground truth template (see Figure 1 in Section I), and this complicated the detection of ground truth templates for all evaluated approaches. Another challenging data set was apache, where only OpenChat achieved the F1-score of 0.714, whereas F1-scores of other approaches remained below 0.5. From the evaluated four methods, Wizardlm2 had the lowest F1-scores, which exceeded 0.5 only in the case of the snmpd data set.
@@
As discussed in Section IV-B, we have used principle P1 to assess the correctness of detected templates. However, not employing this principle would allow us to assess the capabilities of different algorithms to infer ground truth templates from insufficient event log data. In Table IV, relevant results are displayed.
@@
TABLE IV: Performance of Template Detection Methods without P1
@@
OpenChat 19 10 3 5 3 0.559 0.741 0.286 0.714 0.429
Mistral 11 8 - 4 1 0.361 0.727 - 0.615 0.2
Wizardlm2 7 1 - 4 0 0.163 0.091 - 0.615 0
@@
According to Table IV, the performance of all methods deteriorated, and the decline in the F1-score was most notable for Drain. For example, for the sshd data set, the F1-score decreased by 0.5, with the number of correctly detected templates decreasing from 29 to 9. Also, in the case of suricata and apache data sets, the number of correctly identified templates (and thus the F1-score) dropped to 0. Drain achieved its highest F1-score of 0.533 on the snmpd data set.
For LLM-based methods, the decline in the performance of OpenChat was less severe, with its performance not deteriorating on su and suricata data sets. Also, OpenChat achieved the highest F1-score on all five data sets. Mistral managed to reach its best F1-scores of 0.727 and 0.615 on su and snmpd data sets. As for Wizardlm2, which had a fairly modest original performance (see Table III), its F1-score dropped from 0.429 to 0 on the apache data set.
@@
# event log data (two messages in sshd data set) sshd[17250]: Failed none for invalid user edward from 192.168.0.89 port 44460 ssh2 sshd[17266]: Failed none for invalid user edward from 192.168.0.89 port 44632 ssh2 # template detected by Drain <*> Failed none for invalid user edward from 192.168.0.89 port <*> ssh2 # ground truth template (detected by OpenChat and Mistral) sshd[<*>]: Failed <*> for invalid user <*> from <*> port <*> ssh2 # event log data (15 messages in sshd data set) sshd[11356]: Protocol major versions differ for 192.168.3.141: SSH-2.0-OpenSSH_5.3 vs. SSH-9.9-OpenSSH_5.0 sshd[11357]: Protocol major versions differ for 192.168.3.141: SSH-2.0-OpenSSH_5.3 vs. SSH-1.33-OpenSSH_5.0 ... # template detected by Drain <*> Protocol major versions differ for 192.168.3.141: SSH-2.0-OpenSSH_5.3 vs. <*> # template detected by OpenChat sshd[<*>]: Protocol major versions differ for <*>: SSH-<*>-OpenSSH_<*> vs. SSH-<*>-OpenSSH_<*> # ground truth template sshd[<*>]: Protocol major versions differ for <*>: <*> vs. <*>
Figure 4: Inferring ground truth templates from insufficient log data.
@@
That advantage of LLM-based approaches over traditional algorithms is essential if ground truth templates have words that contain delimited-⟨$\langle$*$\rangle$ ∗ ⟩ as a substring, and the template mining algorithm is expected to identify them exactly as they appear in ground truth templates. For example, about 89% of ground truth templates of data sets from Table I have such a nature, and without using the P2 principle for assessing Drain’s performance, the performance would decrease significantly. Also, without using neither P1 nor P2 principle, the performance of OpenChat, Mistral, and Wizardlm2 remained the same as reported in Table IV, because we applied P2 only to Drain during the experiments. On the other hand, disabling both principles reduced the F1-score for Drain to 0 on all data sets, illustrating the strengths of LLM-based approaches over traditional algorithms when complex ground truth templates need to be identified.
As pointed out in a recent paper by Khan et al. [25], existing works have not investigated incorrectly detected templates enough, and studying them helps to understand which types of detection errors are most common for specific algorithms. If $t$ is a template, let $log(t)$ denote the set of all potential messages (i.e., not only the messages in the evaluation data set!) which are matching template $t$. If $log(t_{1})\subset log(t_{2})$, then template $t_{1}$ is called more specific than template $t_{2}$ (i.e., any message matched by $t_{1}$ is also matched by $t_{2}$, but not vice versa). Also, $t_{2}$ is called more general than $t_{1}$. According to [25], an incorrect template is over-generalized (OG) if it is more general than some ground truth template, and is under-generalized (UG) if it is more specific than some ground truth template. If an incorrect template is neither OG nor UG, it is called mixed (MX).
# a ground truth template sshd[<*>]: Server listening on <*> port <*>. # an over-generalized (OG) template sshd[<*>]: <*> # two under-generalized (UG) templates sshd[<*>]: Server listening on 0.0.0.0 port <*>. sshd[<*>]: Server listening on :: port <*>. # a mixed (MX) template <*> Server listening on 0.0.0.0 port <*>
Figure 5: Types of incorrectly detected templates.
Figure 5 provides an example message for each type. For example, the last example from Figure 5 is of type MX since, unlike the ground truth template, it assumes the constant value 0.0.0.0 for the interface IP address field. On the other hand, unlike the ground truth template, it does not assume that the message begins with the program name sshd and ends with the dot character. Table V provides data about the types of incorrect templates from our evaluation.
TABLE V: Incorrectly Detected Templates by Type
@@
As seen in Table V, LLM-based approaches tend to produce very few MX templates, and for OpenChat no such templates were generated on five data sets. In contrast, for Drain, the MX templates formed a significant part among incorrectly detected templates on sshd, su, and suricata data sets. As discussed above, traditional algorithms (including Drain) are not able to identify character patterns inside message words, and our investigation revealed that many incorrect templates from Drain represent the same scenario as illustrated by the MX template from Figure 5.
@@
When manually inspecting OG templates, we discovered that Mistral generated an OG template snmpd[delimited-⟨$\langle$*$\rangle$ ∗ ⟩]: delimited-⟨$\langle$*$\rangle$ ∗ ⟩ on the snmpd data set for all 10 experiment iterations (that template matches all messages in the data set). During some iterations, that template was generated during earlier phases of the event log processing, and since detected templates are used for filtering out log messages for performance optimization reasons (see line 4 in Algorithm 1 in Section III), no further messages were submitted to Mistral for analysis. As can be seen from Table II, that led Mistral to identify only one template snmpd[delimited-⟨$\langle$*$\rangle$ ∗ ⟩]: delimited-⟨$\langle$*$\rangle$ ∗ ⟩ during some experiment iterations.
@@
Since other LLM-based template detection approaches such as LILAC [22] and ULog [23] utilize similar work principles of employing already detected templates as filters, our finding illustrates one design limitation of current LLM-based approaches. Namely, if an underlying LLM is prone to produce overly general templates, a mechanism must be implemented to identify and drop such templates. Unfortunately, whereas some overly general templates like the OG template from Figure 5 are fairly easy to detect, there are other more subtle cases, and their identification is a non-trivial task.
@@
When investigating UG templates, we found that some of them provide novel insights into the event log data despite being incorrect according to ground truth. The most notable case involved the suricata data set, which was the most challenging for all algorithms because Mistral and Wizardlm2 failed to process it, whereas the highest F1-score remained below 0.3 and was achieved by OpenChat (see Table III). In the case of OpenChat, all incorrectly identified templates were of type UG and have been shown in Figure 6. The total number of UG templates was 12, and they were more specific than the ground truth template shown in Figure 1 (Section I), which covered over 99% of the suricata data set.
@@
Also, in Suricata IDS alert messages (see Figure 1 in Section I), the first word of the alert string identifies the signature supplier. For example, in two alerts from Figure 1 the supplier identifiers are ET (free signatures from EmergingThreats) and GPL (community signatures). Another identifier present in the suricata data set is ETPRO (commercial signatures from EmergingThreats).
# template for Suricata IDS alerts from 153 ET signatures suricata[<*>]: [<*>] ET <*> [Classification: <*>] [Priority: <*>] {<*>} <*> -> <*>:<*> # template for Suricata IDS alerts from 30 ETPRO signatures suricata[<*>]: [<*>] ETPRO <*> [Classification: <*>] [Priority: <*>] {<*>} <*> -> <*>:<*> # templates for Suricata IDS alerts from 10 GPL signatures suricata[<*>]: [<*>] GPL DNS named authors attempt [Classification: Attempted Information Leak] [Priority: <*>] {UDP} <*> -> <*>:<*> suricata[<*>]: [<*>] GPL FTP MKD overflow [Classification: Attempted Administrator Privilege Gain] [Priority: <*>] {TCP} <*> -> <*>:<*> ...
Figure 6: Templates from OpenChat describing Suricata IDS alerts.
@@
Despite being incorrectly identified, the UG templates from Figure 6 reveal the fact that IDS alerts have been produced by signatures from three suppliers. Since for ET and ETPRO, a larger number of signatures produced many different alert strings, alerts for both suppliers have been summarized by a single template. However, since for the GPL supplier, only 10 signatures triggered alerts, individual templates have been created for each GPL signature, providing more details for the human analyst. Since OpenChat did not generate any other incorrectly identified templates, templates from Figure 6 can be regarded as a concise and meaningful summary of Suricata IDS alerts, providing useful insights not present in ground truth. Also, it should be pointed out that the output from Figure 6 was not accidentally generated by OpenChat, but exactly the same output was produced during 8 experiment iterations out of 10.
@@
Scenarios from Figure 6 and Figure 4 demonstrate that LLM-based approaches like OpenChat have a potential not only for template detection but also for discovering previously unknown knowledge that is not represented by ground truth. Also, it highlights the fact that detected templates deserve qualitative analysis during evaluations to better assess the knowledge discovery capabilities of different algorithms.
V Discussion
@@
In this section, we will summarize the main findings of the evaluation presented in Section IV.
@@
First, as discussed in Section IV-B, existing performance metrics are not entirely suited for a meaningful comparison of LLM-based and traditional algorithms since the latter have several design limitations. Also, adjusting the ground truth templates according to these limitations (e.g., as proposed in [25]) would not allow us to evaluate the unique capabilities of LLM-based approaches. As a solution, we have proposed two heuristic principles for assessing the correctness of detected templates.
@@
Second, our results show that local LLMs with parameter size 7B can feature a good performance when used without fine-tuning, employing the in-context learning paradigm in an unsupervised fashion. In particular, the LLM-TD algorithm with the OpenChat LLM featured the best performance among all evaluated methods, including Drain, a highly capable traditional algorithm [24]. That allows security analysts to employ local LLMs for unsupervised event log analysis without the need to create labeled data sets. Moreover, since security event logs contain highly sensitive data, they can be analyzed without submitting sensitive information to external service providers.
@@
Whereas existing LLM-based algorithms [22, 23, 20, 21, 19] have attempted to detect one template during each LLM query, the LLM-TD algorithm presented in this paper utilizes LLMs for extracting several templates at once during every query. As the results from Section IV-C indicate, such a detection strategy is a viable alternative to existing approaches.
@@
One drawback of LLM-based template detection algorithms is their need for more computational resources than traditional algorithms require, which leads to longer execution times. For example, during our evaluations, a high-end GPU was used by local LLMs (see Section IV-A), whereas traditional algorithms do not need such dedicated hardware. In addition, Drain featured significantly faster execution times than LLM-TD with OpenChat, Mistral, and Wizardlm2 (see Table II in Section IV-C). Also, in the case of one challenging data set (suricata), Mistral and Wizardlm2 were not able to process it even within 10 hours.
@@
Furthermore, LLM-based approaches have several unique capabilities that traditional algorithms lack. According to our findings (e.g., see Figure 4 in Section IV-C), LLMs can infer informative templates even from insufficient event log data. Also, LLMs are able to detect character patterns inside log message words which can yield novel insights into event log data.
@@
Our study has also identified several design limitations of LLM-based and traditional algorithms, which can interfere with the template detection process. According to recent research literature [20, 26], many traditional algorithms incorrectly assume that log messages matching the same template have the same number of words, and our study has confirmed the negative impact of this assumption. Also, since traditional algorithms consider the log message words as atoms, detected templates are often less informative. As for the limitations of the LLM-based algorithms, they commonly employ already detected templates to exclude log messages from further analysis. However, if overly general templates are identified during the early phases of event log processing, a large fraction of relevant messages could be left unprocessed.
@@
VI Conclusion and Future Work

### v3

@@ 4.3.2 Template Detection Performance
Table 4: Performance of template detection methods (log message batch size 10)
Method Number of correctly detected templates F1-score
(number of all detected templates in parentheses)
sshd su suricata snmpd apache sshd su suricata snmpd apache
OpenChat 25 (31) 10 (13) 3 (15) 7 (7) 5 (7) 0.746 0.769 0.286 1 0.714
Mistral 17 (25) 10 (10) - 0 (1) 0 (2) 0.557 0.87 - 0 0
Wizardlm2 9 (51) 1 (9) - 5 (6) 3 (7) 0.207 0.091 - 0.769 0.429
Drain 29 (44) 9 (16) 3 (80) 6 (8) 2 (4) 0.725 0.621 0.07 0.8 0.364
As for the number of detected templates for 10 experiment iterations, Drain always produced the same result since it is a deterministic algorithm. Also, by its design, Drain assigns all messages in the event log into groups, creating one-message groups for rare messages and thus leaving no messages unprocessed. As mentioned in the previous section, LLM-TD produced the same output during all experiment iterations, behaving similarly to Drain despite the non-deterministic nature of LLMs. However, the templates detected by LLM-TD might not cover the entire event log, and the number of uncovered messages was smallest for OpenChat (four messages for the sshd data set and less for others, see the row for OpenChat with k=10 in Table 2). In contrast, Wizardlm2 produced the largest number of uncovered messages for all data sets (e.g., in the case of sshd and su data sets 1121 and 822, respectively, which complicates their manual analysis).
As discussed in Section 3, one notable difference of LLM-TD from other LLM-based template detection algorithms is its ability to identify multiple templates with one LLM query. This allows to reduce computationally expensive interaction with LLM, and we observed the effect of that property for OpenChat on three data sets. According to Tables 3 and 4, OpenChat was able to detect 25 correct templates with 19 queries on the sshd data set, 10 correct templates with 7 queries on the su data set, and 7 correct templates with 3 queries on the snmpd data set. Mistral featured a similar behavior on the su data set. In contrast, because other LLM-based algorithms can identify at most one template with a single query, the detection of m correct templates will require at least m LLM queries, involving an additional computational overhead.
@@
According to Table 4, OpenChat achieved the best performance on four data sets, whereas Mistral had the highest F1-score on the su data set. The suricata data set was the most challenging for all evaluated approaches since Mistral and Wizardlm2 were not able to process it within 10 hours. For k=2 (the only setting where the processing of the suricata data set completed in 10 hours), the F1-scores of Mistral and Wizardlm2 were 0.25 and 0, respectively. Furthermore, OpenChat and Drain achieved low F1-scores of 0.286 and 0.07 on the suricata data set. The vast majority of the log messages in this data set were Suricata IDS alerts with different numbers of words that correspond to one ground truth template (see Figure 1 in Section 1), and this complicated the detection of ground truth templates for all evaluated approaches. Another challenging data set was apache, where only OpenChat achieved the F1-score of 0.714, whereas F1-scores of other approaches remained below 0.5.
From the evaluated four methods, performance of Mistral and Wizardlm2 was noticeably lower. For example, Wizardlm2 exceeded the F1-score of 0.5 only in the case of the snmpd data set. Also, Mistral yielded the F1-score of 0 on snmpd and apache data sets (the reason was the detection of overly general patterns, an issue which was already mentioned in the previous section and will be discussed in details in Section 4.3.3).
Although k=10 yielded the best F1-scores for the majority of experiment scenarios, we investigated if other values of k would significantly improve the lower F1-scores from Table 4. The F1-score of Mistral increased from 0 to 0.714 on the snmpd data set with k=5, whereas no improvement was observed for the apache data set. As for Wizardlm2, its F1-score increased from 0.091 to 0.4 on the su data set with k=5, and from 0.207 to 0.357 on the sshd data set with k=20. For other scenarios no performance improvements were observed.
@@
As discussed in Section 4.2, we have used principle P1 to assess the correctness of detected templates. However, not employing this principle would allow us to assess the capabilities of different algorithms to infer ground truth templates from insufficient event log data. In Table 5, relevant results are displayed.
@@
Table 5: Performance of template detection methods without P1 (log message batch size 10)
@@
OpenChat 17 10 3 5 3 0.507 0.769 0.286 0.714 0.429
Mistral 11 9 - 0 0 0.361 0.783 - 0 0
Wizardlm2 7 1 - 4 0 0.161 0.091 - 0.615 0
@@
According to Table 5, the performance of all methods deteriorated, and the decline in the F1-score was most notable for Drain. For example, for the sshd data set, the F1-score decreased by 0.5, with the number of correctly detected templates decreasing from 29 to 9. Also, in the case of suricata and apache data sets, the number of correctly identified templates (and thus the F1-score) dropped to 0. Drain achieved its highest F1-score of 0.533 on the snmpd data set.
For LLM-based methods, the decline in the performance of OpenChat was less severe, with its performance not deteriorating on su and suricata data sets. Also, OpenChat achieved the highest F1-score on four data sets. Mistral managed to reach the best F1-score of 0.783 on the su data set, with other F1-scores remaining below 0.4. As for Wizardlm2, its F1-score dropped from 0.429 to 0 on the apache data set, whereas performance degraded less for other data sets.
@@
Figure 4: Inferring ground truth templates from insufficient log data
@@
That advantage of LLM-based approaches over traditional algorithms is essential if ground truth templates have words that contain delimited-⟨$\langle$*$\rangle$ ∗ ⟩ as a substring, and the template mining algorithm is expected to identify them exactly as they appear in ground truth templates. For example, about 89% of ground truth templates of data sets from Table 1 have such a nature, and without using the P2 principle for assessing Drain’s performance, the performance would decrease significantly. Also, without using neither P1 nor P2 principle, the performance of OpenChat, Mistral, and Wizardlm2 remained the same as reported in Table 5, because we applied P2 only to Drain during the experiments. On the other hand, disabling both principles reduced the F1-score for Drain to 0 on all data sets, illustrating the strengths of LLM-based approaches over traditional algorithms when complex ground truth templates need to be identified.
4.3.3 Analysis of Incorrectly Detected Templates
As pointed out in a recent paper by Khan et al. khan , existing works have not investigated incorrectly detected templates enough, and studying them helps to understand which types of detection errors are most common for specific algorithms. If $t$ is a template, let $log(t)$ denote the set of all potential messages (i.e., not only the messages in the evaluation data set!) which are matching template $t$. If $log(t_{1})\subset log(t_{2})$, then template $t_{1}$ is called more specific than template $t_{2}$ (i.e., any message matched by $t_{1}$ is also matched by $t_{2}$, but not vice versa). Also, $t_{2}$ is called more general than $t_{1}$. According to khan , an incorrect template is over-generalized (OG) if it is more general than some ground truth template, and is under-generalized (UG) if it is more specific than some ground truth template. If an incorrect template is neither OG nor UG, it is called mixed (MX).
Figure 5: Types of incorrectly detected templates
Figure 5 provides an example message for each type. For example, the last example from Figure 5 is of type MX since, unlike the ground truth template, it assumes the constant value 0.0.0.0 for the interface IP address field. On the other hand, unlike the ground truth template, it does not assume that the message begins with the program name sshd and ends with the dot character. Table 6 provides data about the types of incorrect templates from our evaluation.
Table 6: Incorrectly detected templates by type (log message batch size 10)
@@ 4.3.3 Analysis of Incorrectly Detected Templates
As seen in Table 6, LLM-based approaches tend to produce very few MX templates, and for OpenChat no such templates were generated on five data sets. In contrast, for Drain, the MX templates formed a significant part among incorrectly detected templates on sshd, su, and suricata data sets. As discussed above, traditional algorithms (including Drain) are not able to identify character patterns inside message words, and our investigation revealed that many incorrect templates from Drain represent the same scenario as illustrated by the MX template from Figure 5. However, on the snmpd and apache data sets, Drain produced just two incorrectly detected templates, and no prominent pattern of MX templates emerged.
@@
When manually inspecting OG templates, we discovered that Mistral generated an OG template snmpd[delimited-⟨$\langle$*$\rangle$ ∗ ⟩]: delimited-⟨$\langle$*$\rangle$ ∗ ⟩ on the snmpd data set during earlier phases of the event log processing. Because detected templates are used for filtering out log messages for performance optimization reasons (see line 4 in Algorithm 1 in Section 3), no further messages were submitted to Mistral for analysis after the detection of this template, although this would have allowed to find additional templates from the remaining event log data. As a result, that led Mistral to identify only one template snmpd[delimited-⟨$\langle$*$\rangle$ ∗ ⟩]: delimited-⟨$\langle$*$\rangle$ ∗ ⟩. On the apache data set, Mistral suffered from the same issue, and the following two OG templates which matched all event log messages prevented the proper analysis of most event log data:
apache2: PHP Notice: delimited-⟨$\langle$*$\rangle$ ∗ ⟩ in delimited-⟨$\langle$*$\rangle$ ∗ ⟩ on line delimited-⟨$\langle$*$\rangle$ ∗ ⟩
apache2: PHP Warning: delimited-⟨$\langle$*$\rangle$ ∗ ⟩ in delimited-⟨$\langle$*$\rangle$ ∗ ⟩ on line delimited-⟨$\langle$*$\rangle$ ∗ ⟩
Also, as discussed in Section 4.3.1, the incorrect exclusion of log messages from processing by OG templates can have side effects on runtime, and allowed Mistral and Wizardlm2 to process the suricata data set with the shortest execution times for k=2.
@@
Since other LLM-based template detection approaches such as LILAC lilac and Lunar ulog utilize similar work principles of employing already detected templates as filters, our finding illustrates one design limitation of current LLM-based approaches. Namely, if an underlying LLM is prone to produce overly general templates, a mechanism must be implemented to identify and drop such templates. Unfortunately, whereas some overly general templates like the OG template from Figure 5 are fairly easy to detect, there are other more subtle cases, and their identification is a non-trivial task. For example, aforementioned two OG templates detected by Mistral on the apache data set contain six words and three delimited-⟨$\langle$*$\rangle$ ∗ ⟩, giving them a look of a regular template and hiding their too general nature.
@@
When investigating UG templates, we found that some of them provide novel insights into the event log data despite being incorrect according to ground truth. The most notable case involved the suricata data set, which was the most challenging for all algorithms because Mistral and Wizardlm2 failed to process it, whereas the highest F1-score remained below 0.3 and was achieved by OpenChat (see Table 4). In the case of OpenChat, all incorrectly identified templates were of type UG and have been shown in Figure 6. The total number of UG templates was 12, and they were more specific than the ground truth template shown in Figure 1 (Section 1), which covered over 99% of the suricata data set.
@@
Also, in Suricata IDS alert messages (see Figure 1 in Section 1), the first word of the alert string identifies the signature supplier. For example, in two alerts from Figure 1 the supplier identifiers are ET (free signatures from EmergingThreats) and GPL (community signatures). Another identifier present in the suricata data set is ETPRO (commercial signatures from EmergingThreats).
Figure 6: Templates from OpenChat describing Suricata IDS alerts
@@
Despite being incorrectly identified, the UG templates from Figure 6 reveal the fact that IDS alerts have been produced by signatures from three suppliers. Since for ET and ETPRO, a larger number of signatures produced many different alert strings, alerts for both suppliers have been summarized by a single template. However, since for the GPL supplier, only 10 signatures triggered alerts, individual templates have been created for each GPL signature, providing more details for the human analyst. Since OpenChat did not generate any other incorrectly identified templates, templates from Figure 6 can be regarded as a concise and meaningful summary of Suricata IDS alerts, providing useful insights not present in ground truth.
@@
Finally, to estimate the capabilities of LLM-TD to process large non-syslog data sets, we applied it to a textual event log of all HTTP requests observed in a large organizational network. The event log contained 15,311,709 messages describing HTTP requests from 36,821 clients, and all messages were following the same format. Figure 7 displays an example anonymized message from this event log (since the log data had a confidential nature, we were not able to release this data set publicly).
Figure 7: Templates from OpenChat describing HTTP requests
When applying LLM-TD with OpenChat to this event log 10 times, it was processed within 6–7 minutes despite its large size of 3.9 GB, producing identical output for all 10 iterations. Furthermore, the memory consumption of LLM-TD remained modest, ranging from 65.5 to 67.4 MB across all iterations (we observed similar memory footprints lower than 70 MB for all other experiments described earlier in this paper).
Although OpenChat was not able to detect a single template for all messages, it was able to meaningfully summarize them with 11 templates (see Figure 7), which represented distinct HTTP methods identified in the event log (for one HTTP method, two templates were created which reflected different usage patterns of this method). However, 17 log messages remained uncovered by templates from Figure 7 and were reported to the end user. When inspecting them more closely, we discovered that six messages represented previously unknown malicious HTTP requests with binary bytes used in the HTTP method field.
@@
Scenarios from Figure 6, Figure 7, and Figure 4 demonstrate that LLM-based approaches like OpenChat have a potential not only for template detection but also for discovering previously unknown knowledge that is not represented by ground truth. Also, it highlights the fact that detected templates deserve qualitative analysis during evaluations to better assess the knowledge discovery capabilities of different algorithms.
5 Discussion
@@
In this section, we will summarize the main findings of the evaluation presented in Section 4.
@@
First, as discussed in Section 4.2, existing performance metrics are not entirely suited for a meaningful comparison of LLM-based and traditional algorithms since the latter have several design limitations. Also, adjusting the ground truth templates according to these limitations (e.g., as proposed in khan ) would not allow us to evaluate the unique capabilities of LLM-based approaches. As a solution, we have proposed two heuristic principles for assessing the correctness of detected templates.
@@
Second, our results show that local LLMs with parameter size 7B can feature a good performance when used without fine-tuning, employing the in-context learning paradigm in an unsupervised fashion. In particular, the LLM-TD algorithm with the OpenChat LLM featured the best performance among all evaluated methods, including Drain, a highly capable traditional algorithm zhu . That allows security analysts to employ local LLMs for unsupervised event log analysis without the need to create labeled data sets. Moreover, since security event logs contain highly sensitive data, they can be analyzed without submitting sensitive information to external service providers.
@@
Whereas existing LLM-based algorithms lilac ; ulog ; divlog ; logppt ; llmparser have attempted to detect one template during each LLM query, the LLM-TD algorithm presented in this paper utilizes LLMs for extracting several templates at once during every query. As the results from Section 4.3 indicate, such a detection strategy is a viable alternative to existing approaches, and it allows to reduce computationally expensive interaction with LLMs.
@@
One drawback of LLM-based template detection algorithms is their need for more computational resources than traditional algorithms require, which leads to longer execution times. For example, during our evaluations, a high-end GPU was used by local LLMs (see Section 4.1), whereas traditional algorithms do not need such dedicated hardware. In addition, Drain featured significantly faster execution times than LLM-TD with OpenChat, Mistral, and Wizardlm2 (see Table 2 in Section 4.3.1). Also, in the case of one challenging data set (suricata), Mistral and Wizardlm2 were not able to process it even within 10 hours for most log message batch sizes.
@@
Furthermore, LLM-based approaches have several unique capabilities that traditional algorithms lack. According to our findings (e.g., see Figure 4 in Section 4.3.2), LLMs can infer informative templates even from insufficient event log data. Also, LLMs are able to detect character patterns inside log message words which can yield novel insights into event log data.
@@
Our study has also identified several design limitations of LLM-based and traditional algorithms, which can interfere with the template detection process. According to recent research literature divlog ; gasimov , many traditional algorithms incorrectly assume that log messages matching the same template have the same number of words, and our study has confirmed the negative impact of this assumption. Also, since traditional algorithms consider the log message words as atoms, detected templates are often less informative. As for the limitations of the LLM-based algorithms, they commonly employ already detected templates to exclude log messages from further analysis. However, if overly general templates are identified during the early phases of event log processing, a large fraction of relevant messages could be left unprocessed.
@@
6 Conclusion and Future Work